At the end of 2014, 5,200,000 Americans were diagnosed with Alzheimer's disease. It is estimated that approximately 200,000 new cases of Alzheimer's are diagnosed each month. What would be the estimated prevalence of Alzheimer's in the United States by the end of March 2015 (for the purpose of this question, assume that no patients die during this period). The 2015 population is estimated to be 320,000,000 people.

18.13 Alzheimer's cases per 1,000 people

12.57 Alzheimer's cases per 1,000 people

16.54 Alzheimer's cases per 1,000 people

14.98 Alzheimer's cases per 1,000 people
Answer #1

We have

Number of Americans were diagnosed with Alzheimer's disease at the end of 2014 = 5200000

Number of Americans were diagnosed with Alzheimer's disease at the end of march 2015 = 5200000 + 3*200000 = 58000000

Population = 320,000, 000

The estimated prevalence of Alzheimer's in the United States by the end of March 2015 = 5200000/320000000

Prevalence = 18.13 Alzheimer's cases per 1000 people.